OpenStudy (anonymous):

solve by square root 2x^2-7=0

OpenStudy (anonymous):

Could you help me with this

OpenStudy (anonymous):

alright, we want to find x so for the current problem:\[2x^2 = 7\]\[x^2 = \frac{7}{2}\]\[x= \pm \sqrt(\frac{7}{2})\] did you understand the steps ? ^_^